Two angles are supplementary. Find their measures if one of them is 20 less than three times the measure of the other.
Answer by nerdybill(7384)

Two angles are supplementary. Find their measures if one of them is 20 less than three times the measure of the other.
.
Let x = one angle
then
180-x = second angle (supplement)
.
180-x = 3x-20
180 = 4x-20
160 = 4x
40 degrees = x
.
second angle:
180 - x = 180-40 = 140 degrees
.
answer: 40 and 140 degrees